Grade 12Electrostatics

An electrostatic field line leaves at angel alpha from point charge +q1 and connects with point charge -q2 at angel beta. Find ration of q1/q2
(Given alpha=60 degree , beta=90 degree)

Aatam
7 Years agoGrade 12

1 Answer

Pooja
7 Years ago
I am unable to attach image with this answer
Hope you understand the solution
Flux coming out from q1 = Flux entering q2
kq1 * 2π{1- cos(alpha/2)} =
                 kq2 * 2π{1- cos(beta/2)}
q* 2sin2(alpha/2)  =  q* 2sin2(beta/2)
q1/q2  =sin2(alpha/2) ÷ sin2(beta/2)
            = sin2(30) ÷ sin2(45) 
            =1/4  ÷  1/2  
            = 1/2